OpenStudy (anonymous):

Two cars, A and B, 100 miles apart, move toward each other. Car A covers 0.7 mile in the first minute, and then moves with an average speed of 0.8 mile per minute. Car B covers 0.6 mile in the first minute, and then moves with an average speed of 0.7 mile per minute. How many miles apart are the cars after 10 minutes? a. 85.2 c. 83.7 b. 14.8 d. 16.3

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[100-\left(\frac{7}{10}+\frac{8}{10}*9\right)-\left(\frac{6}{10}+\frac{7}{10}*9\right)=\frac{426}{5}=85.2 \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your are welcome. Thank you for the medal.